Find the power of i
(-i)42
Possible answers:
A) (-i)2
B) 1
C) i
D) -i
E) -1

Answer by stanbon(75887)

(-i)^42 = (i^3)^(42) = i^(126) = (i^4)^31*i^2 = 1*(-1) = -1
========================
Cheers,
Stan H.
========================
Possible answers:
A) (-i)2
B) 1
C) i
D) -i
E) -1